If 1.47 A of current was passed through an aqueous solution of a metal salt for 525 seconds and 863 mg of metal was deposited, then what was the charge on metal ion?


how would you work this out?


thanks.:)

Grade:12

1 Answers

Pankaj
9 years ago
Using Farady's First Law of Electrolysis.
m = ZIt
or
z = m/It = 863000/ (1.47)(525) =1118.16
Charge on metal ion = 1118.16 x1.60217657 × 10-19coulombs

Thanks & Regards
Pankaj Singh
askIITians Faculty